OpenStudy (anonymous):

if g(x) = x^2 and f(x) = 3x^2-3x-1 then find (g*f) (x)

jigglypuff314 (jigglypuff314):

Hello and Welcome to OpenStudy! :) (g*f)(x) would mean g(x) * f(x) I think so plug in g(x) = x^2 and f(x) = 3x^2-3x-1 and get (x^2)(3x^2 - 3x - 1) = ?

OpenStudy (anonymous):

and that will be the answer?

jigglypuff314 (jigglypuff314):

try distributing the (x^2)(3x^2 - 3x - 1) can you do that? :)

OpenStudy (anonymous):

yeah

jigglypuff314 (jigglypuff314):

tell me what you get then :3

OpenStudy (anonymous):

3x^4-2x^3-1x^2

jigglypuff314 (jigglypuff314):

really close :) try the middle term again (x^2)(-3x) = ?

OpenStudy (anonymous):

oh i am sorry its -2x not -3x

jigglypuff314 (jigglypuff314):

then you are correct :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you